OpenStudy (anonymous):

what is the y-intercept of the graph of 4x+2y=12

OpenStudy (lgbasallote):

change it to the form y = mx + b first

OpenStudy (anonymous):

The y-intercept is the value of the equation when x=0. So, set x=0 and then solve the equation.

OpenStudy (anonymous):

put it in the form y = mx + b -2y = 12 -4x y = -6 +2x, y = 2x - 6 y-intercept (when x=0) y = -6 that is your first point your slope is 2, so when x=1, y= -4 two points define a line. the ordered pairs x y 0 -6 1 -4 2 -2 3 0 source: http://answers.yahoo.com/question/index?qid=20090303155431AAKRS71

OpenStudy (lgbasallote):

that works too :) probably easier

OpenStudy (anonymous):

thank you!

OpenStudy (lgbasallote):

@Rohangrr 2y should not be negtive..

OpenStudy (anonymous):

how do i change it to the y=mx+b form? sorry this is cst review and im currenly talking precal ;~; forgot all about algebra

OpenStudy (lgbasallote):

you have 4x + 2y = 12 to change it to y = mx + b isolate y 2y = 12 - 4x now divide everything by the coefficient of y (which is 2) so y = 6 - 2x rearrange it... y = -2x + 6 the constant is your y-intercept.which in this case is 6